OpenStudy (anonymous):

You live near a bridge that goes over a river. The underside of the bridge is an arch that can be shown by the function y=-0.000371x^2+.827x. the x and y value are measured in feet. How high above the river is the bridge (the top of the arch) and how long is the section of the bridge above the arch? A. bridge is about 460.87 ft. above the river, length is 2229.11ft. B. bridge is about 2229 ft. above the river, length is 460.87 ft. C. bridge is about 460.87 ft. above the river, length is 1114.555 ft. D. bridge is about 230.435 ft. above the river, length is about 22229.11 Please help!!!

OpenStudy (anonymous):

@mathmale will you please help me?

OpenStudy (anonymous):

Calculate the slope of the underside and put it equal to zero (At top the slope should be zero). solve it for x. For this value of x , calculate y. This y is the height of the top.

OpenStudy (anonymous):

Ok how would I do that? Sorry I'm completely lost.

OpenStudy (anonymous):

Ok.First differentiate y with respect to x.

OpenStudy (anonymous):

Umm I'm kinda confused... I don't know what that means.

OpenStudy (anonymous):

So if you have function y = x^n differentiating it with respect to x means y' = n*x^(n-1)

OpenStudy (anonymous):

My computer says its going to restart in a minute so if I don't answer I'll be right back on. Ok so what would I plug into n?

OpenStudy (anonymous):

y=-0.000371x^2+.827x y' = -0.000371*2*x + 0.827*1*x^0 put y' = 0 and solve for x

OpenStudy (anonymous):

Ok I've got that but what do you mean by the *2* and *1* in your equation?

OpenStudy (anonymous):

They are n

OpenStudy (anonymous):

Ok so multiply the .000371 by 2 and the .827 by 1then solve for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I don't know if my post came through but I said it would be 7.42x10^-4x^2+.827x without the x solved for?

OpenStudy (anonymous):

y' = -0.000371*2*x + 0.827*1*x^0 Since y' = 0 -0.000371*2*x + 0.827*1*x^0 = i.e -0.000742x + 0.827 = 0 x = 0.827/0.000742

OpenStudy (anonymous):

Calculate this x and put it in y=-0.000371x^2+.827x to get the height of the top

OpenStudy (anonymous):

Ok thank you I see it now! So x=1114.555256?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the answer would be C right? Since that is the only option? I'm sorry I took so long to get it it's easier for me to get it if it's worked out lol Thanks again!

OpenStudy (anonymous):

No. the answer is not C

OpenStudy (anonymous):

Calculate y for x = 1114.56

OpenStudy (anonymous):

Oh ok I see haha so it's 460.8685984 for y

OpenStudy (anonymous):

Now put y = 0 and solve for x. You should get 2 values of x. subtract them, you'll get length of the section of the bridge above the arch.

OpenStudy (anonymous):

So it's A right?